x2–2x-3 Let f(x) = Find the X-value(s) for which the graph of y = f(x) has a horizontal tangent line. x+2
Answer #1

Aw frus x²2x-3 = 0 (34+2) for horgontal tongent flus (34+2) ene (x? 22-3) —(2122x-3) (+2) (22+2) let inif = fa) = (2+1) (28-2